# QIANets: Quantum-Integrated Adaptive Networks for Reduced Latency and Improved Inference Times in CNN Models

### v1

@@ Abstract
Convolutional neural networks (CNNs) have made significant advances in computer vision tasks, yet their high inference times and latency often limit real-world applicability. While model compression techniques have gained popularity as solutions, they often overlook the critical balance between low latency and uncompromised accuracy. By harnessing quantum-inspired pruning, tensor decomposition and annealing-based matrix factorization – three quantum-inspired concepts – we introduce QIANets: a novel approach of redesigning the traditional GoogLeNet, DenseNet, and ResNet-18 model architectures to process more parameters and computations whilst maintaining low inference times. Despite experimental limitations, the method was tested and evaluated, demonstrating reductions in inference times, along with effective accuracy preservations.
@@ 1 Introduction
The field of computer vision (CV) has recently experienced a substantial rise in interest (Su & Crandall, 2021). This surge has created transformative advancements, driving the development of deep learning models, particularly those based on the convolutional architecture, such as DenseNet, GoogLeNet, and ResNet-18. These methods have significantly optimized neural networks for image processing tasks, achieving state-of-the-art performance across multiple benchmarks. (Anumol, 2023; He et al., 2015; Szegedy et al. 2015) However, the increasing computational complexity, memory consumption, and model size – often comprising millions to billions of parameters – pose substantial challenges for deployment, especially in time-sensitive and computationally-limited scenarios. The demand for low-latency processing in real-time applications, such as image processing and automated CV systems, is critical; compact models are needed for faster responses (Honegger et al., 2014).
To address these issues, researchers have explored various optimization techniques to reduce inference times and latency while maintaining high accuracy. Model compression techniques such as pruning, quantization, and knowledge distillation have shown promise in enhancing model efficiency (Li et al., 2023). Yet, these methods often come with trade-offs that can impact model performance, necessitating a careful balance between energy efficiency and accuracy.
In recent years, the principles of quantum computing have emerged as an avenue for accelerating inference in machine learning (Divya & Dinesh Peter, 2021). Quantum-inspired methods, which leverage phenomena such as quantum optimization algorithms, strive to maintain model performance by reducing computational requirements, thereby offering significant speedups for certain tasks (Pandey et al., 2023). Meanwhile, traditional model compression techniques reduce the size of neural networks by removing less important weights, sacrificing accuracy for lower latency (Francy & Singh, 2024). By integrating concepts from quantum mechanics into convolutional neural network (CNN) models, our approach seeks to address these limitations. We explore the potential of designing CNNs to balance improved inference times with minimal accuracy loss, creating a novel solution.
@@ 2.1 Model Compression Techniques:
Pruning is one of the most effective ways to accelerate CNNs. Cheng et al. (2018) provided a comprehensive review of model compression techniques for deep neural networks (DNNs), specifically focusing on parameter pruning methods: the removal of individual weights based on importance. This significantly reduces model size, while generally preserving the model’s performance.
Despite the advancements in parameter pruning, overall conventional pruning techniques have limitations: 1) when applied during the training phase, as such a method may be costly, and 2) the risk of prematurely removing important data. Hou et al. (2022) proposed a novel methodology called CHEX for training-based channel pruning and regrowing of channels throughout the training process. By employing a column subset selection (CSS) formulation, CHEX allocates and reassigns channels across layers, allowing for significant model compression without requiring a fully pre-trained model.
@@ 2.2 Quantum-Inspired Techniques for CNNs:
Quantum computing is currently recognized as a potential game-changer for various fields, including NLP, due to its ability to process complex data more efficiently than classical computers. Shi et al. (2021) proposed a quantum-inspired architecture for convolutional neural networks (QICNNs), using complex-valued weights to enhance the representational capacity of traditional CNNs. The authors display that their QICNNs achieve higher classification accuracy and faster convergence on datasets compared to standard CNNs. In contrast, our methodology prioritizes structural optimization for greater computational efficiency. We focus on reducing latency and improving inference times by employing various quantum techniques to decrease computational overhead in the tested CNNs.
Hu et al. (2022) set a high standard in the field by addressing the unique challenges of compressing quantum neural networks (QNNs). Their CompVQC framework leverages an alternating direction method of multipliers (ADMM) approach, achieving a remarkable reduction in circuit depth by over 2.5 times with less than 1% accuracy degradation. While their results in QNN compression are impressive, our research introduces a novel first-attempt technique that applies QAOA-inspired pruning, tensor decomposition and quantum annealing-inspired matrix factorization to classical CNNs. Our work can complement their approach, underscoring the potential of integrating quantum concepts to classical networks, and may lead to further improvements in model efficiency.
@@ 3.1 Quantum-inspired pruning
Pruning is a widely applied technique for reducing CNN complexities; early studies have demonstrated the effectiveness of pruning in optimizing neural networks (LeCun et al., 1989; Hanson and Pratt, 1989; Hassibi et al., 1993). Our method implements a new approach to optimization through the Quantum Approximate Optimization Algorithm (QAOA). QAOA is designed to address problems in combinatorial optimization, which involves finding the best solution from a finite set of solutions.
Similarly, we adapt these principles by framing pruning as a probabilistic optimization problem. The goal is to identify the most important weights to retain while allowing others to approach zero. For a neural network layer represented by weights as a tensor: $W\in\mathbb{R}^{C_{out}\times C_{in}\times H\times W}$, we define the importance of each weight using its absolute value: $I_{i,j}=|W_{i,j}|$
To facilitate decision-making regarding weight retention, we normalize these importance scores with the softmax function to derive probabilities:
$P_{i,j}=\sum_{k,l}e^{I_{k,l}}e^{I_{i,j}}$
These probabilities are then utilized in a quantum-inspired decision-making process, where weights are selected for pruning based on a threshold $\mathit{\lambda}$, influenced by a hyperparameter known as layer sparsity.
$R_{i,j}=\begin{cases}1&\text{if }P_{i,j}\geq\lambda\\ 0&\text{otherwise}\end{cases}$
[…]
When implemented, we adopt an iterative approach for pruning the model across multiple stages. Each iteration recalculates the retain mask based on updated probabilities derived from the current weights. To enhance this process, we introduce a neighboring entanglement mechanism: when a weight is pruned, its adjacent weights in the tensor may also be pruned with a specified probability Pentangle. This mechanism simulates quantum entanglement, reflecting the idea that nearby weights may exhibit correlated behavior and can be pruned collectively. For convolutional layers specifically, this sequential pruning strategy is executed over several iterations, progressively reducing the number of parameters in the model while maintaining performance integrity.
@@ 3.2 Tensor Decomposition
Tensor decomposition further reduces the dimensionality of the weight tensor while preserving essential information for accurate predictions. This technique is inspired by Quantum Circuit Learning (QCL), where high-dimensional tensors are decomposed into lower-dimensional forms for efficient training of quantum circuits.
For a weight tensor $W\in\mathbb{R}^{C_{out}\times C_{in}\times H\times W}$, we use Singular Value Decomposition (SVD) on its flattened matrix representation $W_{f}\in\mathbb{R}^{C_{out}\times(C_{in}\cdot H\cdot W)}$, decomposing it as: $W_{f}=U\Sigma V^{T}$
Here, $U\in\mathbb{R}^{C_{out}\times r}$ and $V\in\mathbb{R}^{(C_{in}\cdot H\cdot W)\times r}$ are orthogonal matrices, and $\Sigma\in\mathbb{R}^{r\times r}$ is a diagonal matrix of singular values. The rank r, chosen as a hyperparameter, controls the compression level by retaining only the top r singular values, leading to a reduced rank approximation: $W_{f}\approx U_{r}\Sigma_{r}V_{r}^{T}$. This reduces the number of parameters and computational costs during inference.
After tensor decomposition, the original weight tensor is reconstructed using the truncated matrices, reshaping the compressed weights back into their original form. This process significantly decreases the number of parameters without greatly affecting model performance. We apply this method to each convolutional layer, creating lower-rank approximations to control the model’s capacity.
@@ 3.3 Quantum annealing-inspired matrix factorization
Quantum annealing solves optimization problems by evolving a system toward its lowest energy state. We apply this concept to factorize weight tensors, treating the factorization as an optimization problem aimed at minimizing the difference between the original weights and their factorized representation.
Given a weight matrix $W\in\mathbb{R}^{m\times n}$, we seek to factor it into two lower-dimensional matrices, $W_{1}\in\mathbb{R}^{m\times r}$and $W_{2}\in\mathbb{R}^{r\times n}$, where r is a hyperparameter that controls the rank: $W\approx W_{1}W_{2}$.
The objective is to minimize the reconstruction error: $L(W_{1},W_{2})=\|W-W_{1}W_{2}\|_{F}^{2}$
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm, measuring the difference between the original and factorized matrices. We use an iterative optimization procedure inspired by quantum annealing to minimize loss.
The factorization employs gradient-based optimization, initializing $W_{1}$ and $W_{2}$ randomly. We iteratively minimize the loss function using an optimizer like LBFGS, suitable for small parameter sets and non-convex landscapes. This optimization simulates quantum annealing by gradually reducing the step size, ensuring convergence to a local minimum. Once complete, the compressed weight matrix is defined as: $W_{c}=W_{1}W_{2}$. This compressed matrix replaces the original matrix, reducing model complexity while maintaining performance.
Figure 1: An illustrative diagram showcasing the framework used for Quantum-Inspired Pruning, Tensor Decomposition, and Quantum Annealing-Inspired Matrix Factorization.
@@ 4 Experiments and results
GoogleNet 94% 0.00096 T/I 86% 0.00083 T/I 1:0.52
ResNet 93% 0.000011 T/I 87% 0.00007 T/I 1:0.61
DenseNet 94% 0.000050 T/I 88% 0.00042 T/I 1:0.56
@@ 4.1 Experimental setup
The experiments were conducted within the cloud-based PyTorch framework, utilizing the CIFAR-10 dataset. The CIFAR-10 dataset, which consists of 60,000 32x32 RGB color images in 10 classes, with 6,000 images per class, was split into training and validation sets with an 80/20 ratio. To meet the input size requirements of the models, images were resized to 224x224 pixels. Data preprocessing included normalizing pixel values to the range [-1, 1] using the mean and standard deviation of the dataset. Moreover, data augmentation strategies (random horizontal flipping and random cropping) were applied to generate data variability and improve the models’ performance on unseen data.
All computations were accelerated using CUDA on an NVIDIA A40 GPU via Runpod. Each model underwent training for 50 epochs utilizing the Adam optimizer, with an initial learning rate of 0.001 and weight decay of 1e-4. To run our experiments, the compute used was approximately 1,615,680 TFLOPS-seconds. To ensure consistency across models, batch sizes of 128 were used for training, while batch sizes of 256 were employed for both evaluation and testing within the dataset.
@@ 4.2 Hyperparameter tuning
Hyperparameter tuning is performed using Optuna, a framework that implements a multitude of techniques to optimize certain parameters. Optuna experiments with various combinations of hyperparameters, including batch size, learning rate and ECA Kernel Size, and dynamically adjusts them based on each trial. Following each trial, validation accuracy is calculated to test the effectiveness of current parameters. This information dynamically refines the subsequent parameters, ultimately approaching optimized parameter configurations for model performance.
@@ 4.3 Model specific analysis
To effectively accommodate the unique architectures of each model, we made targeted adjustments to the QIANets method. These modifications were carefully designed to be minimal, ensuring that all models were trained and evaluated under consistent and fair conditions throughout the experiments.
@@ 4.3.1 GoogLeNet
GoogLeNet is a convolutional network with nine multi-scale processing Inception modules. Within our experiments, the QIANets framework targets the these modules, reducing the weight in their convolutions: layer sparsity of 0.1417 (only 14.17% of weights were pruned), while employing a rank of 41 to efficiently decompose and factorize these weights. ECA and Multi-Scale Fusion are applied to the outputs of the modules, integrating multi-scale attention into parallel branches.
Performance Progression Across Epochs: The training process involved 10 trials of 10 epochs each, followed by a final trial of 50 epochs. During Trial 0 of hyperparameter optimization, the model began with a validation accuracy of 21.08% at Epoch 1 and quickly progressed to 70.18% by Epoch 10, indicating rapid learning during the first stages of training. The highest validation accuracy result happened to be 80.19% throughout Trial 5. Ultimately, the final quantum-inspired GoogLeNet’s test accuracy was 86.65%, representing a notable improvement from the earlier 80.19%, and closely approaching the baseline accuracy of 94.29% after fine-tuning.
Loss Reduction: Over the 50 epochs, the model’s loss steadily decreased, showing consistent improvement. It began at 2.5205 in Epoch 1 trial 0 and reduced to 0.8256 by Epoch 50, effectively minimizing error throughout training. A key outcome of this experiment was the final 13.65% reduction in inference time, minimizing to 0.000835 seconds per image, which underscores the efficiency of our approach compared to the baseline GoogLeNet’s 0.000967 seconds.
@@ 4.3.2 DenseNet
We experimented on DenseNet – a CNN structured with 12 dense blocks, with layer-by-layer connections. This intricate connectivity requires careful application of QAOA pruning to ensure that weight removal does not disrupt the model’s residual stream and overall flow of information within the network. To enhance channel-wise interactions, we apply ECA and Multi-Scale Fusion after the dense blocks, allowing the model to leverage both local and global feature representations effectively.
Performance Progression Across Epochs: DenseNet was trained over 10 trials of 10 epochs each, followed by a final trial of 50 epochs to fine-tune performance. During Trial 0, the model began with a validation accuracy of 9.66% at Epoch 1 and exhibited minimal improvement by Epoch 10, reaching 10.34%. In contrast, Trial 1 demonstrated significant learning progression, starting at 27.53% and achieving a remarkable 81.33% by Epoch 10. The highest validation accuracy across all trials peaked at 86.65% in Trial 1, where the model achieved a layer sparsity of approximately 0.3779 (nearly 62% of the weights were pruned while maintaining performance). After extensive fine-tuning, the final quantum-inspired DenseNet achieved a test accuracy of 88.52%, a significant improvement from the earlier 86.65%, and approaching the baseline accuracy of 94.05%.
Loss Reduction: The model demonstrated steady loss reduction throughout the training process, beginning at 2.3028 during Epoch 1 in Trial 0 and decreasing to 0.5606 by Epoch 10 in Trial 1, indicating effective error minimization. This consistent decline reflects the model’s ability to optimize its parameters and improve performance across trials. One of the standout results of this experiment was the final reduction in inference time by 15.20%, dropping to 0.001043 seconds/image, marking a considerable improvement compared to the baseline DenseNet’s 0.00123 seconds.

### v2

@@ Abstract
Convolutional neural networks (CNNs) have made significant advances in computer vision tasks, yet their high inference times and latency often limit real-world applicability. While model compression techniques have gained popularity as solutions, they often overlook the critical balance between low latency and uncompromised accuracy. By harnessing quantum-inspired pruning, tensor decomposition and annealing-based matrix factorization – three quantum-inspired concepts – we introduce QIANets: a novel approach of redesigning the traditional GoogLeNet, DenseNet, and ResNet-18 model architectures to process more parameters and computations whilst maintaining low inference time. Code: https://github.com/quantum-inspired-model-compression
@@ 1 Introduction
The field of computer vision (CV) has recently experienced a substantial rise in interest [1]. This surge has created transformative advancements, driving the development of deep learning models, particularly within convolutional architecture, such as DenseNet [2], GoogLeNet [3], and ResNet-18 [4]. These methods have significantly optimized neural networks for image processing tasks, achieving state-of-the-art performance across multiple benchmarks [5]. However, the increasing computational complexity, memory consumption, and model size (millions to billions of parameters) pose substantial challenges for deployment, especially in time-sensitive and computationally-limited scenarios. The demand for low-latency processing in real-time applications, such as image processing and automated CV systems, is critical; compact models are needed for faster responses [6].
To address these issues, researchers have explored various optimization techniques to reduce inference times and latency while maintaining high accuracy. Model compression techniques such as pruning, quantization, and knowledge distillation have shown promise in enhancing model efficiency [7]. Yet, these methods often come with trade-offs that can impact model performance, necessitating a careful balance between energy efficiency and accuracy.
In recent years, the principles of quantum computing have emerged as an avenue for accelerating inference in machine learning [8]. Quantum-inspired methods, which leverage phenomena such as quantum optimization algorithms, strive to maintain model performance by reducing computational requirements, thereby offering significant speedups for certain tasks [9]. Meanwhile, traditional model compression techniques reduce the size of neural networks by removing less important weights, sacrificing accuracy for lower latency [10]. By integrating concepts from quantum mechanics into convolutional neural network (CNN) models, our approach seeks to address these limitations. We explore the potential of designing CNNs to balance improved inference times with minimal accuracy loss, creating a novel solution.
@@ 2.1 Model Compression Techniques:
Pruning is one of the most effective ways to accelerate CNNs. Cheng et al. (2018) [11] reviewed model compression techniques for deep neural networks (DNNs), focusing on parameter pruning: removing individual weights based on importance to reduce model size while generally preserving performance.
Despite the advancements in parameter pruning, overall conventional pruning techniques have limitations: 1) high cost when applied during training and 2) the risk of prematurely removing important data. Hou et al. (2022) [12] introduced CHEX, a training-based channel pruning and regrowing method that reallocates channels across layers using a column subset selection (CSS) formulation, achieving significant compression without a fully pre-trained model.
@@ 2.2 Quantum-Inspired Techniques for CNNs:
Quantum computing is currently recognized as a potential game-changer for various fields, including NLP, due to its ability to process complex data more efficiently than classical computers.
Shi et al. (2021) [13] proposed a quantum-inspired architecture (QICNNs) with complex-valued weights to enhance CNN representational capacity, achieving higher accuracy and faster convergence on datasets than standard CNNs. In contrast, our methodology prioritizes structural optimization for greater computational efficiency, reducing latency and improving inference times through quantum techniques.
Hu et al. (2022) [14] set a high standard in the field by addressing quantum neural networks (QNNs) compression. Their CompVQC framework leverages an alternating direction method of multipliers (ADMM) approach, achieving a remarkable reduction in circuit depth by over 2.5× with less than 1% accuracy loss. While their results in QNN compression are impressive, our research introduces a novel first-attempt technique that applies QAOA-inspired pruning, tensor decomposition and quantum annealing-inspired matrix factorization to classical CNNs, potentially complementing their approach and enhancing model efficiency.
@@ 3.1 Quantum-Inspired Pruning
We build upon the established technique of pruning to reduce the complexities of CNNs, as demonstrated in early studies ([15]; [16]; [17]). However, we introduce a new optimization way, utilizing the Quantum Approximate Optimization Algorithm (QAOA) [18] to frame pruning as a probabilistic optimization problem. For a neural network layer represented by weights as a tensor: $W\in\mathbb{R}^{C_{out}\times C_{in}\times H\times W}$, we define the importance of each weight using its absolute value:
$I_{i,j}=|W_{i,j}|$ (1)
To facilitate decision-making regarding weight retention, we normalize these importance scores with the softmax function:
$P_{i,j}=\frac{e^{I_{i,j}}}{\sum_{k,l}e^{I_{k,l}}}$ (2)
These probabilities are then used in a quantum-inspired decision-making process. Weights are pruned based on a threshold $\lambda$, influenced by a hyperparameter known as layer sparsity $\alpha$:
$R_{i,j}=\begin{cases}1&\text{if }P_{i,j}\geq\lambda\\ 0&\text{otherwise}\end{cases}$ (3)
Here, $R_{i,j}$ serves as a binary retain mask indicating whether a weight is pruned (set to 0) or retained. The threshold $\lambda$ is calibrated to ensure that approximately $100\alpha\%$ of the weights are pruned.
When implemented, we adopt an iterative approach across multiple stages, recalculating the retain mask based on updated probabilities. To enhance this process, we introduce a neighboring entanglement mechanism: when a weight is pruned, adjacent weights in the tensor may also be pruned with the probability $P_{\text{entangle}}$, simulating quantum entanglement and reflecting correlated behavior among nearby weights. For convolutional layers, this sequential pruning strategy is executed over several iterations, progressively reducing the number of parameters.
@@ 3.2 Tensor Decomposition
Tensor decomposition further reduces the dimensionality of the weight tensor while preserving essential information for accurate predictions. Inspired by Quantum Circuit Learning (QCL), high-dimensional tensors are decomposed into lower-dimensional forms for efficient training of quantum circuits [19].
For a weight tensor $W\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times H\times W}$, we use Singular Value Decomposition (SVD) [20] to its flattened representation $W_{f}\in\mathbb{R}^{C_{\text{out}}\times(C_{\text{in}}\cdot H\cdot W)}$:
$W_{f}=U\Sigma V^{T}$ (4)
Here, $U$ and $V$ are orthogonal matrices, and $\Sigma$ is a diagonal matrix of singular values. The rank $r$, chosen as a hyperparameter, controls the compression by retaining only the top $r$ singular values:
$W_{f}\approx U_{r}\Sigma_{r}V_{r}^{T}$ (5)
After decomposition, we reconstruct the original weight tensor using truncated matrices, significantly decreasing parameters without greatly affecting model performance.
@@ 3.3 Quantum Annealing-Inspired Matrix Factorization
Quantum annealing optimizes systems toward their lowest energy state [21]. We apply this concept to factorize weight tensors, treating it as an optimization problem aimed at minimizing the difference between the original and factorized weights.
Given a weight matrix $W\in\mathbb{R}^{m\times n}$, we factor it into two lower-dimensional matrices, $W_{1}\in\mathbb{R}^{m\times r}$ and $W_{2}\in\mathbb{R}^{r\times n}$, where $r$ is a hyperparameter that controls the rank:
$W\approx W_{1}W_{2}$ (6)
The objective is to minimize the reconstruction error:
$L(W_{1},W_{2})=\|W-W_{1}W_{2}\|_{F}^{2}$ (7)
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm. We use an iterative optimization procedure based on quantum annealing to minimize this loss.
The factorization employs gradient-based optimization, initializing $W_{1}$ and $W_{2}$ randomly. Using an optimizer like LBFGS, suitable for small parameter sets and non-convex landscapes, we simulate quantum annealing by gradually reducing step size to ensure convergence to a local minimum. Once complete, the compressed weight matrix is defined as:
$W_{c}=W_{1}W_{2}$ (8)
This compressed matrix replaces the original matrix, reducing model complexity.
Figure 1: An illustrative diagram showcasing the QIANets framework
@@ 4 Experiments and Results
GoogleNet 94% 0.00096 T/I 86% 0.00083 T/I x1.9
ResNet 93% 0.00011 T/I 87% 0.00007 T/I x1.6
DenseNet 94% 0.000050 T/I 88% 0.000042 T/I x1.8
@@ 4.1 Experimental Setup
The experiments were conducted within the PyTorch framework, utilizing the CIFAR-10 dataset [22]. The CIFAR-10 dataset, which consists of 60,000 32x32 RGB color images in 10 classes, with 6,000 images per class, was split into training and validation sets with an 80/20 ratio. Data preprocessing included resizing images to 224x224 pixels and normalizing pixel values to the range [-1, 1]. Moreover, to generate data variability, data augmentation strategies (random horizontal flipping and cropping) were applied, improving performance on unseen data.
All computations were accelerated using CUDA on an NVIDIA A40 GPU via Runpod. Each model underwent training for 50 epochs utilizing the Adam optimizer (learning rate = 0.001, weight decay = 1e-4), using approximately 1.6 million TFLOPS-seconds of compute. To ensure consistency across models, batch sizes of 128 were used for training, while batch sizes of 256 were employed for both evaluation and testing within the dataset. The training process involved 10 trials of 10 epochs each, followed by a final trial of 50 epochs.
@@ 4.2 Hyperparameter Tuning
Hyperparameter tuning is performed using Optuna, a framework that implements various techniques to optimize certain parameters. Optuna tests various combinations of hyperparameters (batch size, learning rate and ECA Kernel Size) and dynamically adjusts them based on each trial. Following each trial, validation accuracy is calculated to test the effectiveness of current parameters. This information refines the subsequent parameters, ultimately approaching optimized parameter configurations.
@@ 4.3 Model Specific Analysis
To effectively accommodate the unique architectures of each model, we made minimal targeted adjustments to the QIANets method but ensured that all models were trained and evaluated under consistent and fair conditions throughout the experiments. See Appendix A for additional details
@@ 4.3.1 GoogLeNet
GoogLeNet is a convolutional network with nine multi-scale processing Inception modules. The QIANets framework targets these modules, reducing the weight in their convolutions: layer sparsity of 0.1417, while employing a rank of 41 to efficiently decompose and factorize these weights. See Table 2.
Table 2: Training and validation metrics for GoogLeNet
Metric Quantum-Inspired GoogLeNet Baseline GoogLeNet
Training Loss 0.7786 1.7066 (Epoch 1)
Validation Accuracy 80.19% 38.52% (Epoch 1)
Test Loss 0.5732 0.2557
Test Accuracy 86.65% 94.29%
Average Inference Time/Image 0.000835 seconds (13.65% Faster) 0.000967 seconds
@@ 4.3.2 DenseNet
We experimented on DenseNet – a CNN structured with 12 dense blocks, with layer-by-layer connections. This intricate connectivity requires careful application of QAOA pruning to ensure that weight removal does not disrupt the model’s residual stream. See Table 3.
Table 3: Training and validation metrics for DenseNet
Metric Quantum-Inspired DenseNet Baseline DenseNet
Training Loss 0.5351 2.3027
Validation Accuracy 81.33% 10.34%
Test Loss 0.4712 0.2462
Test Accuracy 88.52% 94.05%
Average Inference Time/Image 0.000042 seconds (15.20% faster) 0.000050 seconds